OpenStudy (anonymous):

sin(a) _______ = 1/2 [sin(a+b) + sin(a-b)] a) sin(b) b) cos(a) c) cos(b) d) sin(a)

OpenStudy (anonymous):

@ganeshie8 @hartnn @JackJordan

hartnn (hartnn):

can you expand sin (a+b) and sin(a-b) ? know the formulas?

OpenStudy (anonymous):

no..haha

hartnn (hartnn):

ok, here you go! \(\Large \sin (a+b) = \sin a \cos b +\cos a \sin b\) \(\Large \sin (a-b) = \sin a \cos b -\cos a \sin b\) just plug these in, on the right side.

hartnn (hartnn):

like add them and see what u get...

hartnn (hartnn):

anything getting cancelld out ?

OpenStudy (anonymous):

idk why but i couldn't get anything from it

hartnn (hartnn):

did you add? \(\Large \sin (a+b) + \Large \sin (a-b) \\ \Large = \sin a \cos b +\cos a \sin b+ \sin a \cos b -\cos a \sin b\) now you see any term getting cancelled out ?

OpenStudy (anonymous):

So when you add them, the latter terms cancel. You will be left with two lots of sin(a)cos(b) Now dividing by two, what do we get?

OpenStudy (anonymous):

so is it cos (b)?

hartnn (hartnn):

jack buddy! let him/her do something :P

hartnn (hartnn):

yep, sin a cos b so yeah

OpenStudy (anonymous):

hrtnn why'd you give him the answer? He thought it was cos b

hartnn (hartnn):

we all know who gave out the answer, so cut it out lol